# Benchmarking AI Models in Software Engineering: A Review, Search Tool, and Enhancement Protocol

### v1

@@
Abstract
[…]
Index Terms:
@@ Abstract
I Introduction
@@
A notable example of benchmark limitations in code generation evaluation is HumanEval [1], a widely-used dataset that has been used for assessing large language models (LLMs) such as Codex [1], Gemini [2], and GPT-4 [3]. HumanEval was downloaded 97,44497,44497 , 444 times in February 2025 on a single platform (Hugging Face)  [^1] https://huggingface.co/datasets/openai/openai_humaneval. This high number demonstrates the strong and sustained interest in this benchmark. Despite its broad adoption, HumanEval contains numerous flaws and inconsistencies [4]. For instance, Task 47, which requires computing the median of a numerical list, incorrectly states the median of the list [-10,4,6,1000,10,20] is 151515. When queried on this task, ChatGPT-3.5 Turbo reproduced the same incorrect result (see Figure 1). This suggests potential data contamination and benchmark overfitting which can artificially inflate performance scores.
As HumanEval remains a widely-used benchmark in both AI and software engineering communities, several efforts have sought to expand its language support [5, 6, 7, 8] or improve test coverage [9, 10]. However, these extensions often build upon the original dataset without addressing its fundamental deficiencies, allowing inherent issues—such as flawed canonical solutions, vague problem definitions, incorrect tests, and insufficient coverage—to persist. Moreover, LLM-augmented improvements, such as automatic translation, often lack rigorous quality control as well. Lastly, as models have advanced, HumanEval and similar popular benchmarks have become increasingly saturated, with close to 100% scores for recent models.  [^2] https://paperswithcode.com/sota/code-generation-on-humaneval This necessitates continuous elevation of the complexity of programs to better reflect the capabilities of contemporary models.
With this study we aim to address four key challenges in benchmarking: (1) the fragmentation of benchmark knowledge across tasks, (2) the difficulty of selecting contextually relevant benchmarks, (3) the lack of standardized approaches for benchmark creation and refinement, and (4) existing inherent flaws that limit benchmark utility. To tackle these issues, we systematically map the current benchmark landscape to facilitate informed selection and establish unified guidelines for developing robust and adaptable benchmarks. We further demonstrate the effectiveness of our framework through a case study on the highly-popular benchmark, HumanEval.
More specifically, we first conducted a systematic review of benchmarking efforts in AI4SE from 2014 onward and identified 204 benchmarks from 173 studies. To the best of our knowledge, this is the first comprehensive analysis of AI4SE benchmarks. We then extracted a given benchmark’s key metadata, including objectives, category, programming language, natural language, and relevant tasks, to structure the AI4SE benchmark landscape. Leveraging this data, we developed BenchScout  [^3] Accessible through https://evalpro.online/, an extensible semantic search tool that enables users to efficiently identify relevant benchmarks for specific software engineering tasks. To build BenchScout, we applied clustering techniques to contextual embeddings derived from related studies and benchmark documentation, along with our manually-extracted metadata. To obtain the optimal configuration, we performed a parameter search. Additionally, we employed dimensionality reduction techniques to visualize the AI4SE benchmark landscape. We conducted a user study with 22 participants from both industry and academia to gauge the usability, effectiveness, and intuitiveness of BenchScout. It achieved average scores of 4.5, 4.0, and 4.1 out of 5, respectively.
@@
Figure 1: ChatGPT repeating the same error from HumanEval (Screenshot captured in Dec 2023)
[…]
@@
  • We conducted a comprehensive review of 173 studies, identified 204 AI4SE benchmarks, and analyzed their limitations and gaps (section III),
  • We developed and released BenchScout, an extensible semantic search tool to facilitate locating appropriate AI4SE benchmarks. Our user study with 22 participants demonstrated its effectiveness (section IV),
  • We propose a unified approach, BenchFrame, to improve the quality and reliability of benchmarks. A case study on HumanEval resulted in a refined, peer-reviewed benchmark, HumanEvalNext (section V). We assessed ten recent LLMs on this benchmark and verified significant drops in performance compared to the original HumanEval[1] and the enhanced version, HumanEvalPlus [10],
@@
II Research Questions
We propose a set of research questions (RQ) to guide our study, as presented below.
  • RQ1: What is the current landscape of AI4SE benchmarks?
    1. 1.
      What are the existing benchmarks used in the field of AI4SE?
    2. 2.
      What are the key limitations and shortcomings of commonly used benchmarks?
  • RQ2: How can we effectively process the growing number of benchmarks and extract meaningful insights?
    1. 1.
      How can we design a tool to efficiently parse this dense data and make it accessible to SE researchers and practitioners?
    2. 2.
      How useful is such a tool from users’ perspective?
  • RQ3: How can we develop a general framework to improve AI4SE benchmarks?
    1. 1.
      What aspects of existing benchmarks does this framework address?
    2. 2.
      How would the process of applying this framework to an existing benchmark work?
    3. 3.
      What are the effects of applying this framework to an existing benchmark?
III Existing Benchmarks, a Review
III-A Search Criteria and Quality Assessment
@@
We employed a systematic method to search for, identify, and classify AI4SE benchmarks. This procedure involved three primary stages: conducting structured searches, verifying credibility, and developing a taxonomy. We searched on two platforms, namely Google Scholar and Semantic Scholar using the following keywords: “Benchmarks”, “Software Engineering”, “Large Language Models”, “Evaluation”, and “AI4SE”. We additionally search for benchmarks present in the PapersWithCode datasets collection due to the popularity and wide usage of the platform  [^5] https://paperswithcode.com/datasets.
@@
We selected these keywords based on a preliminary assessment of highly-cited benchmark research. Three authors worked together to revise the search criteria, ensuring they were both relevant and complete. Our selection criteria targeted primary studies published in English from 2014 to 2024. The initial retrieval resulted in 962 papers, which, after removing duplicates, reduced to 240. Two authors reviewed the relevance of the literature. During the quality evaluation, the originality (i.e., status as a primary study), reproducibility, and accessibility of each study were assessed. To capture a thorough set of benchmarks, we also performed forward and backward snowballing. After the manual screening, 155 papers were deemed relevant, with an additional 18 identified through snowballing. Ultimately, the authors worked together to develop and consistently improve a taxonomy to efficiently categorize the benchmarks and extract metadata. Through continuous discussions, we identified initial essential categories while systematically gathering additional details for each study, including DOI and publication date. An iterative strategy was used to design the taxonomy and categorize SE tasks, starting with overarching categories such as reasoning, synthesis, and debugging. When the categories became too broad or lacked cohesion, the authors refined them into more detailed subcategories, establishing a multi-tiered hierarchy.
@@
In our evaluation, we not only considered benchmark datasets from academic sources, but also included those suggested by industry and subsequently adopted by scholars (for example, IBM CODAIT and Aider [11]). Our replication package also serves as a dynamic repository, allowing researchers to contribute additional benchmarks and related papers by submitting a pull request that includes the paper’s DOI.
III-B Results of the Review
Figure 2: Number of published benchmark papers since 2014
@@ III-B Results of the Review
Our review revealed a significant increase in the number of benchmarks published over time (see Figure 2). We identified 204 benchmarks in total, with 47 linked papers published in 2024 alone. Using an exponential projection, we estimate that this number will reach 70 for the year 2025. This highlights the growing impact of AI4SE benchmarking and the need for a comprehensive overview of the existing literature. In the following, we summarize key insights from our review. Due to space constraints we present the created taxonomy along with the benchmarks in each category. For more detailed information about the benchmarks please refer to the appendix and replication package.
TABLE I: Overview of AI4SE benchmarks stemming from HumanEval [1].
Category Name
Original HumanEval [1]
@@
Improved Language Support MultiPL-HumanEval [6]
HumanEval-Fix [7]
HumanEval-Explain [7]
HumanEval-Synthesize [7]
HumanEval-X [8]
Multi-HumanEval [5]
HumanEvalXL [12]
Improved Testing HumanEval+ [10]
HumanEval-MINI [10]
HE-Eval [9]
@@
Instruction-based InstructHumanEval  [^6] https://huggingface.co/datasets/codeparrot/instructhumaneval
Extended EvoEval [13]
@@
HumanEval Benchmark family: Currently, one of the most popular AI4SE benchmarks is HumanEval [1], used to evaluate the performance of many notable code-aware models (e.g., Codex [1], Gemini [2], and GPT-4 [3]). This benchmark is used mainly for code synthesis, though there also exist some variations for code repair and code explanation [7]. Table XVIII presents the family of HumanEval benchmarks. After an in-depth analysis of these benchmarks, we identified the following issues: (1) incorrect tests, (2) lack of proper test coverage, (3) incorrect canonical solutions, and (4) imprecise problem definitions. While there are versions that have improved the language support [5, 6, 7, 8] and test coverage [9, 10], there is no version that contains all the improvements combined nor fixed the original issues. The issues for enhancing the original dataset can be generalized as follows:
@@
TABLE II: Overview of AI4SE benchmarks derived from MBPP [14].
Category Name
Original MBPP [14]
Improved Language Support MultiPL-MBPP [6]
MBXP [5]
Improved Testing MBPP+ [10]
MBPP-Eval [9]
MBPP Benchmark Family: Another AI4SE benchmark, highly similar in style and popularity compared to HumanEval, is MBPP [14]: Mostly Basic Python Problems. It contains nearly a thousand crowdsourced problems, where almost half of it is sanitized and separately released. Furthermore, several enhancements have been published for MBPP (Table XIX). Upon a more in-depth analysis of MBPP and its family of benchmarks, there are many signs suggesting deficient quality. One notable problem is the lack of proper testing, as MBPP originally only has three (rather trivial) tests per problem – which are all revealed in the prompt as well. With such a test suite in place, evaluation metrics become unstable and insignificant for proper comparison. The strength of the written tests and solutions themselves is not only troublesome in the original data but also the sanitized data features many flaws (even in corrected variants [10]). From negligible observations such as poor syntax (e.g., too many spaces, Python method names starting with a capital – this is a common convention to only use for classes) to uncaught bugs and edge cases that break the implementation. While there are enhancements that improve the language support and extend the test cases, they are all built upon inadequate foundations, which renders any MBPP benchmark suboptimal for proper assessment.
@@
TABLE III: Overview of competitive programming, code complexity, and code efficiency benchmarks.
Category Name
Competitive Programming CodeContests [15]
APPS [16]
LiveCodeBench [17]
LeetCode [18]
CodeElo [19]
Code Complexity CoRCoD [20]
GeeksForGeeks (GFG) [21]
CODAIT  [^7] CODAIT-2021 https://ibm.co/4emPBIa
CodeComplex [22]
PythonSaga [23]
Code Efficiency EffiBench [24]
CODAL [25]
PIE [26]
@@
TABLE IV: Overview of data science & domain-specific benchmarks.
Category Benchmark
Data Science DS-1000 [27]
NumpyEval [28]
PandasEval [28]
JuICe [29]
DSP [30]
ExeDS [31]
DSEval [32]
Domain-Specific Bio-Coder [33]
Bio-Coder-Rosalind [33]
WebApp1k [34]
TABLE V: Overview of Mathematical Reasoning Benchmarks.
Category Name
Mathematical Reasoning MATH [35]
MATH500 [36]
MathQA [37]
MathQA-Python [14]
MathQA-X [5]
Līla [38]
MultiArith [39]
GSM8K [40]
GSM-HARD [40]
TheoremQA [41]
PECC [42]
BRIGHT [43]
AMC12888https://huggingface.co/datasets/AI-MO/aimo-validation-amc
TABLE VI: Overview of Natural Language Benchmarks.
Category Name
Text2Code CoNaLa [44]
MCoNaLa [45]
CoNaLa-SO [46]
APPS [16]
APPS-Eval [9]
AixBench [47]
Natural2Code [2]
CoSQA [48]
WebQueryTest [49]
AdvTest [49]
CONCODE [50]
MTPB [51]
CAASD  [52]
Shellcode_IA32  [53]
Odex  [53]
PSB2  [54]
TACO  [55]
Turbulence  [56]

### v2

# Benchmarking AI Models in Software Engineering: A Review, Search Tool, and Unified Approach for Elevating Benchmark Quality

@@
Abstract.
Benchmarks are fundamental for ensuring unified evaluation and reproducibility. The rapid adoption of Artificial Intelligence in Software Engineering (AI4SE) has led to numerous benchmarks for tasks such as code generation and bug repair. However, this proliferation has introduced major challenges: (1) fragmented knowledge of benchmarks across tasks, (2) difficulty in selecting contextually relevant benchmarks, (3) lack of standardization in benchmark creation and enhancement, and finally (4) flaws that limit the utility of benchmarks. Addressing these challenges requires a dual approach: first, systematically mapping the current benchmark ecosystem to enable informed selection, and second, establishing unified guidelines for developing robust and adaptable benchmarks.
In this work, we first provide a comprehensive review of 247 studies from which we identify 273 benchmarks for evaluating AI4SE tasks since 2014. We categorize these benchmarks and analyze their limitations to highlight gaps in current benchmarking practices.
Drawing on the insights derived from the review, we built an extensible semantic search tool, BenchScout, to locate suitable benchmarks. We utilized an automated clustering technique that incorporates embedded contexts of a given benchmark’s associated studies.
We then applied dimensionality reduction and performed a parameter search to automatically optimize the configuration. To evaluate BenchScout’s usability, effectiveness, and intuitiveness, we conducted a user study with 22 participants, yielding average scores of 4.5, 4.0, and 4.1 out of 5, respectively.
Next, to improve benchmarking standards, we propose BenchFrame, a unified approach to elevate benchmarks’ quality. As a case study, we applied BenchFrame to the widely-used HumanEval benchmark to identify and rectify its key shortcomings. This resulted in HumanEvalNext which includes (1) corrected errors, (2) better language conversion, (3) increased test coverage, and (4) increased difficulty. We then assessed 10 state-of-the-art code language models on HumanEval, HumanEvalPlus, and HumanEvalNext. Compared to HumanEval and HumanEvalPlus, models’ performance on HumanEvalNext drops by an average of $31.22\%$ and $19.94\%$ in the pass@1 scores, respectively. This significant drop illustrates the values of continuous improvement for higher quality benchmarks. Moreover, we tackle issues related to the scalability of BenchFrame through a case study involving an agentic pipeline that delivers similar results. Similarly, we confront questions of generalizability by employing BenchFrame on the MBPP dataset, which equally impacts models’ performance comparably.
Lastly, we publicly release the material of our review, user study, and the enhanced benchmark.  [^1] https://github.com/AISE-TUDelft/AI4SE-benchmarks
@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXX
1. Introduction
@@
Figure 1. ChatGPT repeating the same error from HumanEval (Screenshot captured in Dec 2023)
A notable example of benchmark limitations in code generation evaluation is HumanEval (Chen et al., 2021), a widely-used dataset that has been used for assessing large language models (LLMs) such as Codex (Chen et al., 2021), Gemini (Gemini Team, 2023), and GPT-4 (OpenAI, 2023). HumanEval was downloaded 82 thousand times in July 2025 on a single platform (Hugging Face)  [^2] https://huggingface.co/datasets/openai/openai_humaneval, which supports the strong and sustained interest in this benchmark. Despite its broad adoption, HumanEval contains numerous flaws and inconsistencies (van Dam et al., 2024). For instance, Task 47, which requires computing the median of a numerical list, incorrectly states the median of the list [-10,4,6,1000,10,20] is $15$. When queried on this task, ChatGPT-3.5 Turbo reproduced the same incorrect result (see Figure 1). This suggests potential data contamination and benchmark overfitting which can artificially inflate performance scores.
As HumanEval remains a widely-used benchmark in both AI and software engineering communities, several efforts have sought to expand its language support (Athiwaratkun et al., 2022; Cassano et al., 2022; Muennighoff et al., 2023; Zheng et al., 2023) or improve test coverage (Dong et al., 2023; Liu et al., 2023a). However, these extensions often build upon the original dataset without addressing its fundamental deficiencies, allowing inherent issues—such as flawed canonical solutions, vague problem definitions, incorrect tests, and insufficient coverage—to persist. Moreover, LLM-augmented improvements, such as automatic translation, often lack rigorous quality control as well. Lastly, as models have advanced, HumanEval and similar popular benchmarks have become increasingly saturated, with close to 100% scores for recent models.  [^3] https://evalplus.github.io/leaderboard.html This necessitates continuous elevation of the complexity of programs to better reflect the capabilities of contemporary models.
Our central objective is to address the systemic challenges within AI4SE benchmarking. we aim to address four key challenges in benchmarking: (1) the fragmentation of benchmark knowledge across tasks, (2) the difficulty of selecting contextually relevant benchmarks, (3) the lack of standardized approaches for benchmark creation and refinement, and (4) existing inherent flaws that limit benchmark utility. To address this, we use a three-stage approach where each stage builds upon the findings of the previous one.
More specifically, we first conducted a systematic review of benchmarking efforts in AI4SE from 2014 onward and identified 273 benchmarks from 247 studies. We then extracted a given benchmark’s key metadata, including (1) objectives, (2) category, (3) programming language, (4) natural language, (5) relevant tasks, (6) extent to which tests are present, (7) scale of the dataset, (8) dataset source, (9) language specificity, (10) maintenance adequacy, (11) whether the dataset is reviewed or not, (12) whether it is frequently used or not, (13) the licensing, and (14) how it was created, to structure the AI4SE benchmark landscape. The findings from our review confirm a fragmented and difficult-to-navigate ecosystem, which motivated us to use these data to develop tooling to easily navigate the landscape of AI4SE benchmarking. BenchScout  [^4] Accessible through https://evalpro.online/ is an extensible semantic search tool that enables users to efficiently identify relevant benchmarks for specific software engineering tasks. To build BenchScout, we applied clustering techniques to contextual embeddings derived from related studies and benchmark documentation, along with our manually-extracted metadata. Additionally, we employed dimensionality reduction techniques to visualize the AI4SE benchmark landscape. We conducted a user study with 22 participants from both industry and academia to gauge the usability, effectiveness, and intuitiveness of BenchScout. It achieved average scores of 4.5, 4.0, and 4.1 out of 5, respectively.
While BenchScout aids navigation, our analysis confirms that many foundational benchmarks are flawed, which means a systematic solution is required to deal with them. Based on the identified gaps and limitations in current benchmarks, we introduce BenchFrame, a unified approach to enhance the quality of both existing and new benchmarks. To demonstrate its efficacy, we conduct a case study on HumanEval and present HumanEvalNext as an enhanced version.
When evaluating performance using HumanEval (original) and HumanEvalNext (our improved version based on the BenchFrame), we observe a substantial decline in pass@1 accuracy. Across ten state-of-the-art open-source code models, the average pass@1 score decreases by 31.2%, with a median drop of 26.0%. Performance remains significantly lower even on HumanEvalPlus (Liu et al., 2023a), an enhanced version of HumanEval, with an average decline of $19.94\%$ in pass@1 scores. These results highlight the importance of continuously refining benchmarks to better guide future research and provide more realistic assessments of model performance.
@@
  • We conducted a comprehensive review of 247 studies, identified 273 AI4SE benchmarks, and analyzed their limitations and gaps (Section 3),
  • We developed and released BenchScout, an extensible semantic search tool to facilitate locating appropriate AI4SE benchmarks. Our user study with 22 participants demonstrated its effectiveness (Section 4),
  • We propose a unified approach, BenchFrame, to improve the quality and reliability of benchmarks. A case study on HumanEval resulted in a refined, peer-reviewed benchmark, HumanEvalNext (Section 5). We assessed ten recent LLMs on this benchmark and verified significant drops in performance compared to the original HumanEval(Chen et al., 2021) and the enhanced version, HumanEvalPlus (Liu et al., 2023a),
@@
2. Research Questions
To systematically address the challenges outlined in the introduction, we structure our study around our central research objective: to diagnose the fragmentation and quality issues within the AI4SE benchmark ecosystem, provide practical tools to navigate this complexity, and establish a rigorous methodology for future quality improvement. We deconstruct this objective into three guiding Research Questions (RQs), which are designed to be addressed sequentially.
  • RQ1: Diagnosing the problem: To what extent is the AI4SE benchmarking landscape fragmented, and what are its primary, systemic limitations?
  • RQ2: Providing a Navigational Aid: How can a semantic search and visualization tool improve the discoverability and selection of relevant AI4SE benchmarks?
  • RQ3: Proposing a Systemic Solution: What constitutes a rigorous and scalable methodology for improving benchmark quality, and what is its measurable impact on model evaluation?
These research questions directly map to the three-stage approach of our study. RQ1 fulfills the diagnostic part of our objective by mapping the landscape and identifying its core problems. The findings from RQ1 directly motivate RQ2, which provides a practical tool to help the community navigate the benchmark fragmentation we uncovered. Finally, RQ3 addresses the root cause of the quality issues found in RQ1 by proposing and validating a systematic framework for improvement, thereby fulfilling the final part of our research objective.
3. Existing Benchmarks, a Review
3.1. Search Criteria and Quality Assessment
@@
We employed a systematic method to search for, identify, and classify AI4SE benchmarks. This procedure involved three primary stages: conducting structured searches, verifying credibility, and developing a taxonomy. We searched on two platforms, namely Google Scholar and Semantic Scholar using the following keywords: “Benchmarks”, “Software Engineering”, “Large Language Models”, “Evaluation”, and “AI4SE”. We additionally search for benchmarks present in the PapersWithCode datasets collection due to the popularity and wide usage of the platform  [^6] This platform has since been discontinued and no longer accessible. The previously used url was https://paperswithcode.com/datasets.
We selected these keywords based on a preliminary assessment of highly-cited benchmark research. Three authors worked together to revise the search criteria to ensure they were both relevant and complete. Our selection criteria targeted primary studies published in English from 2014 to 2025. After the initial paper collection process, duplicates were removed. Two authors reviewed the relevant of the identified pieces of literature, during which the originality (that is, the status as primary study), reproducibility, and accessibility of each study was accessed.
More specifically, for originality, we accessed whether a study was a primary study that introduced a new benchmark or a significant extension of an existing one. This filter was set in place to distinguish between benchmark-proposing papers from studies that merely applied them with a certain level preprocessing. in terms of Reproducibility and Accessibility, we evaluated whether the benchmark and its associated resources were made publicly available. We looked for the presence of a public repository (e.g., GitHub, HuggingFace, etc.) which would allow the community to access and use the benchmark. After this phase, we performed forward and backward snowballing. A total of 247 papers were found through this process.
@@
After the papers were collected, the authors worked together to develop and consistently improve a taxonomy to efficiently categorize the benchmarks and extract metadata. Through continuous discussions, we identified initial essential categories while systematically gathering additional details for each study, including DOI and publication date. An iterative strategy was used to design the taxonomy and categorize SE tasks, starting with overarching categories such as reasoning, synthesis, and debugging. When the categories became too broad or lacked cohesion, the authors refined them into more detailed subcategories, establishing a multi-tiered hierarchy.
@@
In our evaluation, we not only considered benchmark datasets from academic sources, but also included those suggested by industry and subsequently adopted by scholars (for example, IBM CODAIT and Aider (CODAIT, 2021)). Our replication package also serves as a dynamic repository, allowing researchers to contribute additional benchmarks and related papers by submitting a pull request that includes the paper’s DOI.
After finalizing the taxonomy, and the list of benchmarks, we extracted (1) objectives, (2) categories, (3) programming language, (4) natural languages, (5) relevant tasks, (6) extent to which tests are present, (7) scale of the dataset, (8) dataset source, (9) language specificity, (10) maintenance adequacy, (11) whether the dataset is reviewed or not, (12) whether it is frequently used or not, (13) the licensing, and (14) how it was created, for each of the benchmarks.
3.2. Results of the Review
Figure 2. Number of published benchmark papers in the last decade. (Jan 2014, Aug 2025)
@@ 3.2. Results of the Review
Our review revealed a significant increase in the number of benchmarks published over time (see Figure 2). We identified 273 benchmarks in total, with 71 published in 2024 alone. Using an exponential projection, we estimate that this number will reach 109 for the year 2025. This highlights the growing impact of AI4SE benchmarking and the need for a comprehensive overview of the existing literature.
Additionally, upon analyzing the distribution of the tasks included in the review, we find that a great portion of benchmarks are related to Code Generation, more specifically, 31.8% of all the benchmarks included. Along with Project-Level Code Generation/Analysis and Code Understanding (with respectively 18.7% and 12.0%), these make up more than 60% of all benchmarks included. This, in itself, underscores a lack of attention to tasks such as vulnerability detection and code retrieval/search (cf. Figure 3).
The data reveal a notable change over time in the distribution of research categories throughout the observed period (cf. Figure 4). A key insight is the shift from code repair and maintenance to automated synthesis, marked by two opposing trends. Initially dominant, the Repair & Maintenance category gradually experiences a notable decline in its share. In contrast, starting around 2020, the Code Generation category gains prominence and reaches its peak in 2022, which coincides with the advancement of large-scale language models. In the recent years presented, there is an emerging trend of thematic diversification; as the focus on Code Generation recedes from its high, categories like Code Understanding and Retrieval & Search show a relative resurgence. This suggests a community-wide transition towards tackling additional challenges related to verification, integration, and the understanding of algorithmically generated outputs.
In the following, we summarize key insights from our review.
Code Generation (31.8%) Project-Level Analysis (18.7%) Code Understanding (12.0%) Math & Data (10.7%) Repair & Maintenance (10.7%) Other (6.7%) Vulnerability Detection (5.4%) Retrieval & Search (4.0%)
Figure 3. The distribution of the high-level categories of the benchmarks included in the study.
Figure 4. The progression of task distribution for the benchmarks included in the study;
Table 1. Overview of AI4SE benchmarks stemming from HumanEval (Chen et al., 2021).
Category Name Language(s) # Tests
Original HumanEval (Chen et al., 2021) Python Avg. 7.7
@@
Improved Language Support MultiPL-HumanEval (Cassano et al., 2022) 18 programming languages Avg. 7.7
HumanEval-Fix (Muennighoff et al., 2023) 6 programming languages Avg. 7.7
HumanEval-Explain (Muennighoff et al., 2023) 6 programming languages Avg. 7.7
HumanEval-Synthesize (Muennighoff et al., 2023) 6 programming languages Avg. 7.7
HumanEval-X (Zheng et al., 2023) 5 programming languages Avg. 7.7
Multi-HumanEval (Athiwaratkun et al., 2022) 12 programming languages Avg. 7.7
HumanEvalXL (Peng et al., 2024) 12PLs, 23NLs Avg. 8.33
mHumanEval (Raihan et al., 2024) Python, 204NLs Avg. 7.7
Improved Testing HumanEval+ (Liu et al., 2023a) Python Scaled ×80
HumanEval-MINI (Liu et al., 2023a) Python Scaled ×47
HE-Eval (Dong et al., 2023) Python Scaled ×14
@@
Instruction-based InstructHumanEval  [^7] https://huggingface.co/datasets/codeparrot/instructhumaneval Python Avg. 7.7
Extended EvoEval (Xia et al., 2024) Python Multiple categories,
scaled with EvalPlus
@@
HumanEval Benchmark family: Currently, one of the most popular AI4SE benchmarks is HumanEval (Chen et al., 2021), used to evaluate the performance of many notable code-aware models (e.g., Codex (Chen et al., 2021), Gemini (Gemini Team, 2023), and GPT-4 (OpenAI, 2023)). This benchmark is used mainly for code synthesis, though there also exist some variations for code repair and code explanation (Muennighoff et al., 2023). Table 1 presents the family of HumanEval benchmarks. After an in-depth analysis of these benchmarks, we identified the following issues: (1) incorrect tests, (2) lack of proper test coverage, (3) incorrect canonical solutions, and (4) imprecise problem definitions. While there are versions that have improved the language support (Athiwaratkun et al., 2022; Cassano et al., 2022; Muennighoff et al., 2023; Zheng et al., 2023) and test coverage (Dong et al., 2023; Liu et al., 2023a), there is no version that contains all the improvements combined nor fixed the original issues. The issues for enhancing the original dataset can be generalized as follows:
@@
Table 2. Overview of AI4SE benchmarks derived from MBPP (Austin et al., 2021).
Category Name Language(s) # Problems
Original MBPP (Austin et al., 2021) Python 974
Improved Language Support MultiPL-MBPP (Cassano et al., 2022) 18 programming languages 354-397 per language
MBXP (Athiwaratkun et al., 2022) 13 programming languages 848-974 per language
Improved Testing MBPP+ (Liu et al., 2023a) Python 427
MBPP-Eval (Dong et al., 2023) Python 974
MBPP Benchmark Family: Another AI4SE benchmark, highly similar in style and popularity compared to HumanEval, is MBPP (Austin et al., 2021): Mostly Basic Python Problems. It contains nearly a thousand crowdsourced problems, where almost half of it is sanitized and separately released. Furthermore, several enhancements have been published for MBPP (Table 2). Upon a more in-depth analysis of MBPP and its family of benchmarks, there are many signs suggesting deficient quality. One notable problem is the lack of proper testing, as MBPP originally only has three (rather trivial) tests per problem – which are all revealed in the prompt as well. With such a test suite in place, evaluation metrics become unstable and insignificant for proper comparison. The strength of the written tests and solutions themselves is not only troublesome in the original data but also the sanitized data features many flaws (even in corrected variants (Liu et al., 2023a)). From negligible observations such as poor syntax (e.g., too many spaces, Python method names starting with a capital – this is a common convention to only use for classes) to uncaught bugs and edge cases that break the implementation. While there are enhancements that improve the language support and extend the test cases, they are all built upon inadequate foundations, which renders any MBPP benchmark suboptimal for proper assessment.
@@
Table 3. Overview of competitive programming, code complexity, and code efficiency benchmarks.
Category Name Language(s) # Tests
Competitive Programming CodeContests (Li et al., 2022a) 12 programming languages Avg. 203.7
APPS (Hendrycks et al., 2021a) Python Avg. 13.2
LiveCodeBench (Jain et al., 2024) Python Avg. 17.23
LeetCode (Tian et al., 2023) Python Avg. 135
CodeElo (Quan et al., 2025) N/A 408 problems
FVAPPS (Dougherty and Mehta, 2025) Python 4715 problems
KodCode (Xu et al., 2025b) 12 programming languages Avg. 7.52
FC2Code (Liu et al., 2022) Python 320 flowcharts
CodeContests+ (Wang et al., 2025b) Python 11,690 problems
LiveCodeBenchPro (Zheng et al., 2025a) Python 584 problems
Code Complexity CoRCoD (Sikka et al., 2019) Java 932
GeeksForGeeks (GFG) (Moudgalya et al., 2023) C++, Python ±1,400 per lang.&categ
CODAIT  [^8] CODAIT-2021 https://ibm.co/4emPBIa Python 4,000,000
CodeComplex (Baik et al., 2024) Java, Python 4,900 per language
PythonSaga (Yadav and Singh, 2024) Python 185
Code Efficiency EffiBench (Huang et al., 2024) Python Self-defined, avg. 100
CODAL (Weyssow et al., 2024) Python 3 ref. / problem
PIE (Shypula et al., 2024) C++ 82.5(median, train)
COFFE (Peng et al., 2025) Python 756
@@
Table 4. Overview of data science & domain-specific benchmarks. (JN refers to Jupyter Notebooks.)
Name Language(s) # Tests Comment
DS-1000 (Lai et al., 2022) Python Avg. 1.6 7 DS/ML libraries
NumpyEval (Zan et al., 2022b) Python Avg. 20 functions NumPy (101 problems)
(Avg. 1 variable)
PandasEval (Zan et al., 2022b) Python Avg. 20 functions Pandas (101 problems)
(Avg. 1 variable)
JuICe (Agashe et al., 2019) Python, JN N/A Cell completion
(1.5M/3.7K train test)
DSP (Chandel et al., 2022) Python, JN Available Cell completion
(1,119 problems)
ExeDS (Huang et al., 2022) Python, JN Execution Based Cell generation
(ground truth), 534 tasks
DSEval (Zhang et al., 2024a) Python custom appraoch Models Evaluated via the DSEval
Approach from the Paper
TorchDataEval (Zan et al., 2022a) Python 50 Private PyTorch data library
MonkeyEval (Zan et al., 2022a) Python 101 Private Pandas library fork
BeatNumEval (Zan et al., 2022a) Python 101 Private NumPy variant
Bio-Coder (Tang et al., 2024b) Python 1,026 Identify and import necessary
Java 1,243 classes for given task
Bio-Coder-Rosalind (Tang et al., 2024b) Python 253 golden solution Generate code for question
WebApp1k (Cui, 2024) React Available evaluates whether a model can
generate React web-app
Table 5. Overview of Mathematical Reasoning Benchmarks.
Name Language(s) # Problems
MATH (Hendrycks et al., 2021b) English 12,500
MATH500 (Lightman et al., 2023) English 500
MathQA (Amini et al., 2019) English 37,297
MathQA-Python (Austin et al., 2021) Python 23,914
MathQA-X (Athiwaratkun et al., 2022) Python, Java, JS 1,883 per language
LiLa(Mishra et al., 2022) Python 133,815 questions, 358,769 programs
MultiArith (Roy and Roth, 2015) English 600
GSM8K (Gao et al., 2022) English 1,320
GSM-HARD (Gao et al., 2022) English 1,320
TheoremQA (Chen et al., 2023) English 800
PECC (Haller et al., 2024) Python 1,006
BRIGHT (Su et al., 2024) English 395
AMC12999https://huggingface.co/datasets/AI-MO/aimo-validation-amc English 82
Table 6. Overview of Natural Language Benchmarks.
Category Name Language(s) No. of Problems
Text2Code CoNaLa (Yin et al., 2018) English $\rightarrow$ Python 2,879
MCoNaLa (Wang et al., 2022) {Spanish, Japanese, 896
Russian} $\rightarrow$ Python
CoNaLa-SO (Orlanski and Gittens, 2021) English $\rightarrow$ Python 10,000
APPS (Hendrycks et al., 2021a) English $\rightarrow$ Python 10,000
APPS-Eval (Dong et al., 2023) English $\rightarrow$ Python 10,000
AixBench (Hao et al., 2022) English, Chinese $\rightarrow$ Java 175
Natural2Code (Gemini Team, 2023) English $\rightarrow$ Python Unknown
CoSQA (Huang et al., 2021) English $\rightarrow$ Python 20,604
WebQueryTest (Lu et al., 2021) English $\rightarrow$ Python 1,046
AdvTest (Lu et al., 2021) English $\rightarrow$ Python 280,634
CONCODE (Iyer et al., 2018) English $\rightarrow$ Java 104,000
MTPB (Nijkamp et al., 2023) English $\rightarrow$ Python 115
CAASD  (Zhang et al., 2024c) English $\rightarrow$ Python 72
Shellcode_IA32  (Liguori et al., 2022) English $\rightarrow$ IA32/Shell 3200
Odex  (Liguori et al., 2022) {Spanish, Japanese, 945 {90, 164,
Russian, English} 252, 439}
$\qquad\qquad\qquad\rightarrow$ Python 1707 test total
PSB2  (Helmuth et al., 2021) English $\rightarrow$ {Clojure, Python} 25
question-answer pairs
TACO  (Li et al., 2023a) English $\rightarrow$ Python 1,539,152 on 26,433 distinct tasks
Turbulence  (Honarvar et al., 2023) English $\rightarrow$ Python 60 (with 420 total test cases)